OpenStudy (anonymous):

I will award a medal to whoever helps me solve the following system of equations. 3y=9 9y+2x=-12

OpenStudy (compassionate):

Hi, can you write your system to be more clear?

OpenStudy (anonymous):

all I have is 3y=9 9y+2x=-12

OpenStudy (compassionate):

3y = 9 9y + 2x = -12 Is that correct?

OpenStudy (anonymous):

Yes

OpenStudy (compassionate):

Solve for y. y = 3 plug it into the second equation. 9y + 2x = -12 9(3) + 2x = -12 27 + 2x = -12 2x = -39 x = 19.5 Plug 19.5 in for x. 9y + 2(19.5) = -12 9y = 39 = -12 9y = -51 y = 5.7 (19.5, 5.7)

OpenStudy (anonymous):

Didnt help

OpenStudy (compassionate):

That's it. I am right. You did not give me the correct systems, as I mentioned. 3y = 9 isn't even a proper format for systems.

OpenStudy (compassionate):

The variables to solve for must be a..z, or A..Z.

OpenStudy (anonymous):

my answers are: 1/3, 15/2), (-1/3, -15/2), (8,8) or (-39/2, 3)

OpenStudy (compassionate):

Hi, If you convert (19.5, 5.7) into a fraction, you'll get: \[(\frac{ 39 }{ 2}, 3)\]

OpenStudy (anonymous):

still confused

OpenStudy (compassionate):

Take. The. Decimals. And. Convert. Them. To. Fractions.

OpenStudy (compassionate):

I just converted them for you, dear.